OpenStudy (anonymous):

D and E are the midpoints of the sides of triangle ABC. If DE=2x, BC=x+10, DE=?

OpenStudy (anonymous):

PLEASE HELP :(((

OpenStudy (anonymous):

 \[DE=\frac{ 1 }{2 }BC,BC=2DE,x+10=2(2x)=4x\] calculate x and then DE=2x

OpenStudy (anonymous):

How will you solve for x? :(

OpenStudy (anonymous):

4x-x=10 3x=10 x=10/3 DE=2x=2*10/3=20/3